The perimeter of a rectangle is 48 m. The length is 3 m less than twice the width. Find the length and the width.
Answer by TimothyLamb(4379)

2L + 2W = 48
L = 2W - 3
---
2L + 2W = 48
2(2W - 3) + 2W = 48
4W - 6 + 2W = 48
6W = 48+6
W = (48+6)/6
W = 9
L = 2W - 3
L = 2*9 - 3
L = 15
---
answer:
W = 9 m
L = 15 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php